# A mathematical theory of topological invariants of quantum spin systems

### v1

@@ 1 Introduction
The study of gapped phases of quantum matter at zero temperature is an important area of theoretical physics. Much conceptual progress has been made by assuming that gapped phases can be described by topological quantum field theory (TQFT). For example, the celebrated Quantum Hall Effect is captured by Chern-Simons field theory. However, the precise relation between gapped phases of matter and TQFTs is not understood. Recently, new mathematically rigorous approaches to classifying gapped phases of matter having been developed (see [1, 2, 3, 4] for the case of one-dimensional systems, [5, 6, 7, 8, 9, 10] for the case of two-dimensional systems, and [11] for systems in an arbitrary number of dimensions). They enable one to assign indices to gapped states of infinite-volume quantum systems invariant under symmetries. The main property of these indices, also referred to as topological invariants, is that they do not vary along suitably-defined continuous paths in the space of states. In some cases, the indices can be related to physical quantities, such as the zero-temperature Hall conductance, thereby explaining the robustness of the latter.
@@
The primary goal of this paper is provide a proper mathematical framework for the constructions of [7, 11, 8] and to interpret topological invariants of gapped states as lattice analogs of ’t Hooft anomalies in Quantum Field Theory. The secondary goal is to generalize the construction in various directions. In particular, we show how to define topological invariants of lattice spin systems living on well-behaved subsets of the lattice. This generalization makes explicit that the invariants take values in a vector space which is determined by the asymptotic geometry of the subset.
@@
While our work concerns quantum lattice spin systems, we take inspiration from Quantum Field Theory (QFT). These two subjects are connected via the bulk-boundary correspondence. One aspect of this conjectural correspondence is that topological invariants of gapped states with symmetries are related to ’t Hooft anomalies of symmetries of the boundary field theory. [^1] This assumes, of course, that a field-theoretic description of boundary modes exists, which is far from obvious. It is usually said that ’t Hooft anomalies are obstructions to gauging a global symmetry of a QFT [12]. A possible mathematical interpretation of this ”physical definition” is that an ’t Hooft anomaly is an obstruction to defining a local action of the group of gauge transformations on the algebra of local observables of a QFT. If this interpretation is correct, then the presence of ’t Hooft anomaly is a purely kinematic statement which involves neither the Hamiltonian nor the vacuum state of the field theory. It is not clear if conventional markers of ’t Hooft anomalies, such as anomalous Ward identities for vacuum correlators of currents, are implied by a kinematic statement. Proving or disproving this is currently out of reach because of gaps in the mathematical foundations of QFT. The mathematical theory of quantum lattice systems, on the other hand, is sufficiently mature and enables us to address the problem of ’t Hooft anomalies from the bulk side of the bulk-boundary correspondence. In this paper we show that topological invariants of gapped states of lattice systems, such as the zero-temperature Hall conductance, can be interpreted as obstructions to promoting a symmetry of a gapped state to a gauge symmetry. Dynamics enters this statement only though the state.
@@
The main novelty of the paper is a new formulation of locality on the lattice. Building on ideas introduced in [7, 11, 8], we define for any (possibly unbounded) region in the lattice a space of derivations that are approximately localized on that region. For sufficiently regular regions, we show that these spaces behave as expected under natural operations like the commutator. To combine them into single geometric object we show that they form a cosheaf on a certain site, ie. a category with a Groethendieck topology. The utility of Groethendieck topologies in describing spaces of functions with a prescribed asymptotic behavior is well-known to analysts, see e.g. [13, 14], and here we apply the same idea in the non-commutative context. We call the resulting global geometric object a local Lie algebra. An onsite action of a Lie group yields a representation of a certain local Lie algebra by derivations, and we show that the invariants defined in [11] can be phrased as purely algebraic invariants of this representation. Namely, an obstruction exists to defining a representation which acts by state-preserving derivations, and this obstruction lives in the homology of a certain DGLA associated naturally to the local Lie algebra of state-preserving derivations. The invariants defined in [11] then arise from a natural pairing of this homology with the Čech cohomology of the sphere at infinity. Aside from clarifying their mathematical meaning, this also shows that the invariants defined in [11] do not depend on certain choices present in their construction.
@@
The content of the paper is as follows. In Section 2 we axiomatize the notion of an infinitesimal local symmetry by defining local Lie algebras abstractly in terms of cosheaves on a site. We also introduce the Čech functor which assigns a DGLAs to a local Lie algebra, and will serve as the main link between lattice geometry and algebraic topology. In Section 3 we turn to infinitesimal symmetries on the lattice, introducing the space of derivations almost-localized on a region, and prove the main properties of these spaces, some of which hold only for sufficiently regular regions. In section 4 we identify a suitable category of such regular regions in ${\mathbb{R}}^{n}$, the category ${\mathcal{C}\mathcal{S}}_{n}$ of fuzzy semilinear sets. We identify a Groethendieck topology on ${\mathcal{C}\mathcal{S}}_{n}$, show that infinitesimal symmetries of any gapped state $\psi$ of a quantum lattice system on ${\mathbb{R}}^{n}$ can be described by a local Lie algebra ${\mathfrak{D}}^{\psi}_{al}$ over ${\mathcal{C}\mathcal{S}}_{n}$. We also study gapped states invariant under an action of a compact Lie group $G$ and define a $G$-equivariant analog of ${\mathfrak{D}}^{\psi}_{al}$. In Section 6 we construct invariants of $G$-invariant gapped states. The construction is along the lines of [11, 8] and uses an inhomogeneous Maurer-Cartan equation. We show that these invariants are obstruction for promoting the symmetry $G$ to a local symmetry of the state $\psi$. We also explain how to generalize the construction of invariants to lattice systems living on sufficiently nice (asymptotically conical) subsets of ${\mathbb{R}}^{n}$ and show that their invariants take values in a space which depends on the asymptotic geometry of the subset. This goes beyond what one can access using the TQFT heuristics. In appendix A we isolate some proofs from section 3, and Appendix B develops some properties of the inhomogeneous Maurer-Cartan equation.
@@
We are grateful to Owen Gwilliam and Ezra Getzler for discussions. The work of A. A. and A. K. was supported by the Simons Investigator Award. B. Y. would like to thank Yu-An Chen and Peking University for hospitality. A. K. would like to thank Yau Mathematical Sciences Center, Tsinghua University, for the same.
@@ 2.1 Locality and (pre-)cosheaves
Let $M$ be a manifold and $Open(M)$ be the category whose objects are open subsets of $M$ and the set of morphisms from an open $U$ to an open $V$ is the singleton or the empty set depending on whether $U\subseteq V$ or $U\nsubseteq V$. Composition of morphisms is uniquely defined. A pre-cosheaf ${\mathfrak{F}}$ on $M$ with values in a category ${\mathcal{C}}$ is a functor ${\mathfrak{F}}:Open(M)\rightarrow{\mathcal{C}}$. Thus for every inclusion of opens $U\subseteq V$ one is given a co-restriction morphism $e_{VU}:{\mathfrak{F}}(U)\rightarrow{\mathfrak{F}}(V)$ such that for any three opens $U\subseteq V\subseteq W$ one has $e_{WV}\circ e_{VU}=e_{WU}$. Pre-cosheaves (as well as pre-sheaves, which are functors from the opposite category of $Open(M)$ to ${\mathcal{C}}$) can be used to describe local data on $M$. This form of locality is rather weak, since it does not require ${\mathfrak{F}}(U\cup V)$ to be expressible through ${\mathfrak{F}}(U)$ and ${\mathfrak{F}}(V)$.
As an example, consider the Lie algebra of gauge transformations, i.e. the Lie algebra ${\mathfrak{G}}(M):=C^{\infty}(M,{\mathfrak{g}})$ of smooth functions on $M$ with values in a finite-dimensional Lie algebra ${\mathfrak{g}}$. It is a global object attached to $M$. To ”localize” it, for any open $U\subseteq M$ we define the Lie algebra ${\mathfrak{G}}(U)$ to be the space of smooth ${\mathfrak{g}}$-valued functions on $M$ whose closed support is contained in $U$. In particular, ${\mathfrak{G}}(\varnothing)=0$. For any inclusion of opens $U\subseteq V$ we have an injective homomorphism of Lie algebras $\iota_{VU}:{\mathfrak{G}}(U)\rightarrow{\mathfrak{G}}(V)$ such that the Lie algebras ${\mathfrak{G}}(U)$ assemble into a pre-cosheaf ${\mathfrak{G}}$ of Lie algebras on $M$. The injectivity of the maps $\iota_{VU}$ for all $U\subseteq V$ means that the pre-cosheaf is coflasque. [^2] The terminology comes from sheaf theory, where a pre-sheaf ${\mathcal{F}}:Open(M)^{opp}\rightarrow{\mathcal{C}}$ is called flasque if for any $U\subseteq V$ the restriction morphism ${\mathcal{F}}(V)\rightarrow{\mathcal{F}}(U)$ is a surjection.
@@
A local Lie algebra over $(X,\leq)$ is a coflasque pre-cosheaf of Lie algebras with Property I which is also a cosheaf of vector spaces over the corresponding site. A morphism of local Lie algebras is a morphism of the underlying pre-cosheaf of Lie algebras.
@@
Let ${\mathfrak{F}}$ be a local Lie algebra. Then for any $U\leq V$ the Lie algebra ${\mathfrak{F}}(U)$ is an ideal in ${\mathfrak{F}}(V)$. One can equivalently define a local Lie algebra as a coflasque pre-cosheaf of Lie algebras over $(X,\leq)$ which is a cosheaf of vector spaces and such that all co-restriction maps are inclusions of Lie ideals.
@@
In this paper all Lie algebras will be Fréchet-Lie algebras and all morphisms will be continuous. We define a local Fréchet-Lie algebra over $(X,\leq)$ to be a coflasque pre-cosheaf of Fréchet-Lie algebras with Property I which is also a cosheaf of vector spaces. It might have been more natural to impose the cosheaf condition in the category of Fréchet spaces. This is less restrictive than the cosheaf condition in the category of vector spaces.
@@ 2.2 DGLA attached to a cover
with the Čech differential $\partial:C_{n+1}\rightarrow C_{n}$. Here some linear order on $I$ has been chosen.
@@
See [16], Corollary 4.3. Note that [16] uses the term ”flabby” instead of ”coflasque”. ∎
Let $C^{aug}({\mathfrak{U}},W;{\mathcal{F}})=\{C_{\bullet}({\mathfrak{U}},W;{\mathcal{F}})\rightarrow{\mathcal{F}}(A)\}$ be the augmented Čech complex.
@@
Let ${\mathfrak{U}}$ be a cover of $W$ indexed by $I$. Let $V$ be a vector space with a basis $e_{i}$, $i\in I$ and let $f^{i},i\in I$, be the dual basis of $V^{*}$. The 1-shifted augmented Čech complex of ${\mathfrak{F}}$ w. r. to ${\mathfrak{U}}$ is naturally identified with a sub-complex of the DGLA $({\mathfrak{F}}(W)\otimes\Lambda^{\bullet}V,\partial)$, where $\partial$ is contraction with $\sum_{i}f^{i}$. It is easy to check that this sub-complex is closed w. r. to the graded Lie bracket thanks to Property I. By Lemma 2.1, the resulting DGLA is acyclic. ∎
@@
Functoriality in ${\mathfrak{F}}$ is clear. Functoriality in ${\mathfrak{U}}$ is written Čech component-wise for ${\mathsf{p}}^{\mathfrak{V}}\in{\mathfrak{M}}({\mathfrak{V}})$
@@
for $i_{0}<i_{2}<\cdots<i_{k}$. Here $k$ denotes the Čech grading. ∎
We will call this functor the Čech functor. We will need some variants of the Čech functor. First, we can define a graded local Lie algebra over a distributive pre-ordered set $(X,\leq)$ in an obvious manner. The construction of the acyclic DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{F}})$ works in this case as well, except that it may have components in negative degrees.
Second, we define a pointed DGLA as a DGLA equipped with a distinguished central cycle of degree $-2$ (”curvature”). A morphism of pointed DGLAs is a DGLA morphism which preserves the distinguished central cycle. [^5] The category of pointed DGLAs as defined here is a full subcategory of the category of curved DGLAs as defined in [17]. For a curved DGLA with curvature ${\mathsf{B}}$, ${\mathsf{B}}$ is not required to be central and the derivation $\partial$ satisfies $\partial^{2}={\rm ad}_{\mathsf{B}}$. We say that a graded local Lie algebra ${\mathfrak{F}}$ over $(X,\leq)$ with a terminal object $T\in X$ is pointed if it is equipped with a distinguished central element ${\mathsf{B}}\in{\mathfrak{F}}(T)$ of degree $-2$. Morphisms in the category of pointed graded local Lie algebras are required to preserve the distinguished element. Then the DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},T;{\mathfrak{F}})$ is an acyclic pointed DGLA, the distinguished central cycle being ${\mathsf{B}}\in C^{aug}_{-1}({\mathfrak{U}},T;{\mathfrak{F}})$. Of course, since the DGLA is acyclic, this central cycle is exact. The construction of a pointed DGLA from a pointed graded local Lie algebra and a cover of $T$ is functorial in both arguments.

### v2

@@ 1 Introduction
The study of gapped phases of quantum matter at zero temperature is an important area of theoretical physics. Much conceptual progress has been made by assuming that gapped phases can be described by topological quantum field theory (TQFT). For example, the celebrated Quantum Hall Effect is captured by Chern-Simons field theory. However, the precise relation between gapped phases of matter and TQFTs is not understood. Recently, new mathematically rigorous approaches to classifying gapped phases of matter have been developed (see [1, 2, 3, 4] for the case of one-dimensional systems, [5, 6, 7, 8, 9, 10] for the case of two-dimensional systems, and [11] for systems in an arbitrary number of dimensions). They enable one to assign indices to gapped states of infinite-volume quantum systems invariant under symmetries. The main property of these indices, also referred to as topological invariants, is that they do not vary along suitably-defined continuous paths in the space of states. In some cases, the indices can be related to physical quantities, such as the zero-temperature Hall conductance, thereby explaining the robustness of the latter.
@@
The primary goal of this paper is provide a proper mathematical framework for the constructions of [11, 8] and to interpret topological invariants of gapped states as lattice analogs of ’t Hooft anomalies in Quantum Field Theory. The secondary goal is to generalize the construction in various directions. In particular, we show how to define topological invariants of lattice spin systems confined to well-behaved subsets of the lattice. This generalization makes explicit that the invariants take values in a vector space which is determined by the asymptotic geometry of the subset.
@@
While our work concerns quantum lattice systems, we take inspiration from Quantum Field Theory (QFT). These two subjects are connected via the bulk-boundary correspondence. One aspect of this conjectural correspondence is that topological invariants of gapped states with symmetries are related to ’t Hooft anomalies of symmetries of the boundary field theory. [^1] This assumes, of course, that a field-theoretic description of boundary degrees of freedom exists, which is far from obvious. It is usually said that ’t Hooft anomalies are obstructions to gauging a global symmetry of a QFT [12]. A possible mathematical interpretation of this statement is that an ’t Hooft anomaly is an obstruction to defining a local action of the group of gauge transformations on the algebra of local observables of a QFT. Assuming this interpretation, the presence of an ’t Hooft anomaly is a purely kinematic statement which involves neither the Hamiltonian nor the vacuum state of the field theory. It is not clear if conventional markers of ’t Hooft anomalies, such as anomalous Ward identities for vacuum correlators of currents, are implied by a kinematic statement. Proving or disproving this is currently out of reach because of gaps in the mathematical foundations of QFT. The mathematical theory of quantum lattice systems, on the other hand, is sufficiently mature and enables us to address the problem of ’t Hooft anomalies from the bulk side of the bulk-boundary correspondence. In this paper we show that topological invariants of gapped states of lattice systems, such as the zero-temperature Hall conductance, can be interpreted as obstructions to promoting a symmetry of a gapped state to a gauge symmetry. Dynamics enters this statement only though the state.
@@
The main novelty of the paper is a new formulation of locality on a lattice. Building on ideas introduced in [7, 11, 8], we define for any (possibly unbounded) region of the lattice a space of derivations that are approximately localized on that region. For sufficiently regular regions, we show that these spaces behave as expected under natural operations like the commutator. To combine them into a single geometric object we show that they form a cosheaf on a certain site, i. e. a category with a Grothendieck topology. The utility of Grothendieck topologies in describing spaces of functions with a prescribed asymptotic behavior is well-known to analysts, see e.g. [13, 14], and here we apply the same idea in a non-commutative context. We call the resulting global geometric object a local Lie algebra. An on-site action of a Lie group on a lattice system yields a representation of a certain local Lie algebra by derivations, and we show that the invariants defined in [11] can be phrased as purely algebraic invariants of this representation. Namely, an obstruction exists to defining a representation which acts by state-preserving derivations, and this obstruction takes value in the homology of a certain DGLA associated naturally to the local Lie algebra of state-preserving derivations. The invariants defined in [11] then arise from a natural pairing of this homology with the Čech cohomology of the sphere at infinity. Aside from clarifying their mathematical meaning, this also shows that the invariants defined in [11] do not depend on certain choices present in their construction.
@@
The content of the paper is as follows. In Section 2 we axiomatize the notion of an infinitesimal local symmetry by defining local Lie algebras abstractly in terms of cosheaves on a site. We also introduce the Čech functor which assigns a DGLAs to a local Lie algebra, and will serve as the main link between lattice geometry and algebraic topology. In Section 3 we turn to infinitesimal symmetries on the lattice, introducing the space of derivations approximately localized on a region, and prove the main properties of these spaces, some of which hold only for sufficiently regular regions. In section 4 we identify a suitable category of such regular regions in ${\mathbb{R}}^{n}$, the category ${\mathcal{C}\mathcal{S}}_{n}$ of fuzzy semilinear sets. We identify a Grothendieck topology on ${\mathcal{C}\mathcal{S}}_{n}$ and show that infinitesimal symmetries of any gapped state $\psi$ of a quantum lattice system on ${\mathbb{R}}^{n}$ can be described by a local Lie algebra ${\mathfrak{D}}^{\psi}_{al}$ over ${\mathcal{C}\mathcal{S}}_{n}$. We also study gapped states invariant under an action of a compact Lie group $G$ and define a $G$-equivariant analog of ${\mathfrak{D}}^{\psi}_{al}$. In Section 6 we construct invariants of $G$-invariant gapped states. The construction is along the lines of [11, 8] and uses an inhomogeneous Maurer-Cartan equation. We show that these invariants are obstructions for promoting the symmetry $G$ to a local symmetry of the state $\psi$. We also explain how to generalize the construction of invariants to lattice systems defined on sufficiently nice (asymptotically conical) subsets of ${\mathbb{R}}^{n}$ and show that their invariants take values in a space which depends on the asymptotic geometry of the subset. This goes beyond what one can access using the TQFT heuristics. In Appendix A we isolate some proofs necessary for Section 3, and Appendix B develops some properties of the inhomogeneous Maurer-Cartan equation.
@@
We are grateful to Owen Gwilliam, Ezra Getzler, and Bas Janssens for discussions. The work of A. A. and A. K. was supported by the Simons Investigator Award. B. Y. would like to thank Yu-An Chen and Peking University for hospitality during the final stages of this work. A. K. would like to thank Yau Mathematical Sciences Center, Tsinghua University, for the same.
@@ 2.1 Locality and (pre-)cosheaves
Let $M$ be a manifold and $Open(M)$ be the category whose objects are open subsets of $M$, and the set of morphisms from an open $U$ to an open $V$ is the singleton or the empty set depending on whether $U\subseteq V$ or $U\nsubseteq V$. Composition of morphisms is uniquely defined. A pre-cosheaf ${\mathfrak{F}}$ on $M$ with values in a category ${\mathcal{C}}$ is a functor ${\mathfrak{F}}:Open(M)\rightarrow{\mathcal{C}}$. Thus for every inclusion of opens $U\subseteq V$ one is given a co-restriction morphism $e_{VU}:{\mathfrak{F}}(U)\rightarrow{\mathfrak{F}}(V)$ such that for any three opens $U\subseteq V\subseteq W$ one has $e_{WV}\circ e_{VU}=e_{WU}$. Pre-cosheaves (as well as pre-sheaves, which are functors from the opposite category of $Open(M)$ to ${\mathcal{C}}$) can be used to describe local data on $M$. This form of locality is rather weak, since it does not require ${\mathfrak{F}}(U\cup V)$ to be expressible through ${\mathfrak{F}}(U)$ and ${\mathfrak{F}}(V)$.
As an example, consider the Lie algebra of gauge transformations, i.e. the Lie algebra ${\mathfrak{G}}(M):=C^{\infty}(M,{\mathfrak{g}})$ of smooth functions on $M$ with values in a finite-dimensional Lie algebra ${\mathfrak{g}}$. It is a global object attached to $M$. To “localize” it, for any open $U\subseteq M$ we define the Lie algebra ${\mathfrak{G}}(U)$ to be the space of smooth ${\mathfrak{g}}$-valued functions on $M$ whose closed support is contained in $U$. In particular, ${\mathfrak{G}}(\varnothing)=0$. For any inclusion of opens $U\subseteq V$ we have a homomorphism of Lie algebras $\iota_{VU}:{\mathfrak{G}}(U)\rightarrow{\mathfrak{G}}(V)$ such that the Lie algebras ${\mathfrak{G}}(U)$ assemble into a pre-cosheaf ${\mathfrak{G}}$ of Lie algebras on $M$. This is a coflasque pre-cosheaf, ie. all its structure maps $\iota_{VU}$ are injective. [^2] The terminology comes from sheaf theory, where a pre-sheaf ${\mathcal{F}}:Open(M)^{opp}\rightarrow{\mathcal{C}}$ is called flasque if for any $U\subseteq V$ the restriction morphism ${\mathcal{F}}(V)\rightarrow{\mathcal{F}}(U)$ is a surjection.
@@
A local Lie algebra over $(X,\leq)$ is a coflasque pre-cosheaf of Lie algebras with Property I which is also a cosheaf of vector spaces over the corresponding site. A morphism of local Lie algebras is a morphism of the underlying pre-cosheaves of Lie algebras.
@@
Let ${\mathfrak{F}}$ be a local Lie algebra. Then for any $U\leq V$ the Lie algebra ${\mathfrak{F}}(U)$ is an ideal in ${\mathfrak{F}}(V)$. One can equivalently define a local Lie algebra as a pre-cosheaf of Lie algebras over $(X,\leq)$ which is a cosheaf of vector spaces and such that all co-restriction maps are inclusions of Lie ideals.
@@
In this paper all Lie algebras will be Fréchet-Lie algebras and all morphisms will be continuous. We define a local Fréchet-Lie algebra over $(X,\leq)$ to be a coflasque pre-cosheaf of Fréchet-Lie algebras with Property I which is also a cosheaf of vector spaces.
The following Lemma will be useful in future sections to check the cosheaf property:
Lemma 2.1.
Let ${\mathfrak{F}}$ be a pre-cosheaf of vector spaces on a distributive lattice $X$, with co-restriction maps $\iota_{U,V}:U\to V$. Then ${\mathfrak{F}}$ is a cosheaf on $X$ (with the topology of finite covers) iff for any $U,V\in X$ the following sequence is exact
[…]
Proof.
We must show that for any $U\in X$ and every covering $U_{1}\vee\ldots\vee U_{n}=U$, the sequence of vector spaces
[…]
is exact, where $\alpha=\sum_{i<j}\iota_{U_{i}\wedge U_{j},U_{i}}-\iota_{U_{i}\wedge U_{j},U_{j}}$ and $\beta=\sum_{i}\iota_{U_{i},U}$. Exactness of the last three terms follows from an easy induction and the associativity of the join operation. For exactness of the first three terms of the sequence, we also proceed by induction: suppose the result holds for all covers of cardinality $n-1$, and let $U_{1},\ldots,U_{n}\in L$ be a cover of $U$. Suppose $(s_{1},\ldots,s_{n})\in\bigoplus_{i=1}^{n}{\mathfrak{F}}(U_{i})$ lies in the kernel of $\beta$. Let $V:=\bigvee_{i=1}^{n-1}U_{i}$. An application of (3) with $U_{n}$ and $V$ shows that $s_{n}=\iota_{U_{n}\cap V,U}(t)$ for some $t\in V\wedge U_{n}=(U_{1}\wedge U_{n})\vee...\vee(U_{n-1}\vee U_{n})$, and by right-exactness of (4) this shows that $s_{n}=\sum_{i=1}^{n-1}\iota_{U_{i}\cap U_{n},U_{n}}(v_{i})$ for some $v_{i}\in{\mathfrak{F}}(U_{i})$. Finally we write
$(s_{1},\ldots,s_{n})=(s_{1}+w_{1},\ldots,s_{n-1}+w_{n-1},0)+(-w_{1},\ldots,-w_{n-1},s_{n})$
@@ 2.2 DGLA attached to a cover
(where some linear order on $I$ has been chosen) with the differential is the Čech differential $\partial:C_{n+1}\rightarrow C_{n}$ given by $\partial=\sum_{j=0}^{n}(-1)^{j}\lambda_{j}$, where $\lambda_{j}$ is the canonical map
${\mathfrak{F}}(U_{i_{0}}\wedge\ldots\wedge U_{i_{n}})\to{\mathfrak{F}}(U_{i_{0}}\wedge\ldots\widehat{U_{i_{j}}}\ldots\wedge U_{i_{n}}).$
@@
See [16], Corollary 4.3. Note that [16] uses the term “flabby” instead of “coflasque”. ∎
Let $C^{aug}({\mathfrak{U}},W;{\mathfrak{F}})=\{C_{\bullet}({\mathfrak{U}},W;{\mathfrak{F}})\rightarrow{\mathfrak{F}}(W)\}$ be the augmented Čech complex.
@@
Let ${\mathfrak{U}}$ be a cover of $W$ indexed by $I$. Let $V$ be a vector space with a basis $e_{i}$, $i\in I$ and let $f^{i},i\in I$, be the dual basis of $V^{*}$. The 1-shifted augmented Čech complex of ${\mathfrak{F}}$ with respect to ${\mathfrak{U}}$ is naturally identified with a sub-complex of the DGLA $({\mathfrak{F}}(W)\otimes\Lambda^{\bullet}V,\partial)$, where $\partial$ is contraction with $\sum_{i}f^{i}$. It is easy to check that this sub-complex is closed with respect to the graded Lie bracket thanks to Property I. By Lemma 2.2, the resulting DGLA is acyclic. ∎
@@
Functoriality in ${\mathfrak{F}}$ is clear. Functoriality in ${\mathfrak{U}}$ is written Čech component-wise for ${\mathsf{p}}^{\mathfrak{V}}\in C_{k}({\mathfrak{V}},W;{\mathfrak{F}})$:
@@
for $i_{0}<i_{2}<\cdots<i_{k}$. ∎
We will call $C^{aug}_{\bullet+1}(-,W;-)$ the Čech functor. We will need some variants of the Čech functor. First, we can define a graded local Lie algebra over a distributive pre-ordered set $(X,\leq)$ in an obvious manner. The construction of the acyclic DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{F}})$ works in this case as well, except that it may have components in negative degrees.
Second, we define a pointed DGLA as a DGLA equipped with a distinguished central cycle of degree $-2$ (which we call the curvature). A morphism of pointed DGLAs is a DGLA morphism which preserves the distinguished central cycle. [^5] The category of pointed DGLAs as defined here is a full subcategory of the category of curved DGLAs as defined in [17]. There, for a curved DGLA with curvature ${\mathsf{B}}$, ${\mathsf{B}}$ is not required to be central and the derivation $\partial$ satisfies $\partial^{2}={\rm ad}_{\mathsf{B}}$. We say that a graded local Lie algebra ${\mathfrak{F}}$ over $(X,\leq)$ with a terminal object $T\in X$ is pointed if it is equipped with a distinguished central element ${\mathsf{B}}\in{\mathfrak{F}}(T)$ of degree $-2$. Morphisms in the category of pointed graded local Lie algebras are required to preserve the distinguished element. Then the DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},T;{\mathfrak{F}})$ is an acyclic pointed DGLA, the distinguished central cycle being ${\mathsf{B}}\in C^{aug}_{-1}({\mathfrak{U}},T;{\mathfrak{F}})$. Of course, since the DGLA is acyclic, this central cycle is exact. The construction of a pointed DGLA from a pointed graded local Lie algebra and a cover of $T$ is functorial in both arguments.